# Weight-Space Linear Recurrent Neural Networks

### v1

@@ C.4 Dynamical System Reconstruction
The LV experiment is performed for 1500 epochs with a batch size of 1024. Training is conducted with a teacher forcing probability $p_{\text{forcing}}=1.0$, meaning the model is always fed the ground truth inputs during training. This is because this is a memorization task, and the goal is for the model to predict the next token knowing the previous one. LSTM and GRU use the same hyperparameters, except with hidden states of sizes 2450 and 2850 respectively (see Table 8).
@@
C.5 Time Series Classification
For time series classification tasks, encompassing both the UEA datasets and the Spirals dataset, models are consistently trained in the non-AR mode, with the categorical cross-entropy loss. Across all these classification experiments, root weight are evolved without weight clipping, and no dynamic tanh activation is applied to their final outputs. Key training hyperparameters exhibit some variation across these diverse datasets: the learning rate is $10^{-5}$ for the Spirals dataset and most UEA datasets (e.g., Ethanol, Heartbeat, Motor, SCP1, SCP2), with the Worms dataset being an exception at $10^{-6}$. The number of training epochs varies widely, ranging from 800 for the Worms dataset, 4000 for Spirals, up to 6500 for the Ethanol UEA dataset, with other UEA datasets generally trained for several thousand epochs. Given our limitation of 16GB available VRAM memory, batch sizes also differ significantly; for instance, the Worms dataset uses a batch size of 40, other UEA datasets use batch sizes typically in the hundreds (from approximately 280 to 560), and the Spirals dataset employs a large batch size of 10000. Regarding data preprocessing, input data normalization is applied for several UEA datasets (specifically Ethanol, Heartbeat, SCP1, and SCP2), but it is not used for others like EigenWorms and MotorImagery, nor is it required for the Spirals dataset.
Appendix D Additional Results
D.1 2D Image Experiments
@@ Appendix D Additional Results
D.2 Spirals & Neural CDEs
Table 10 reveals several limitations of WARP on the toy Spirals dataset originally introduced to test Neural CDEs Kidger [2022]. We find that at the same parameter count, WARP not only struggles to achieve 100% accuracy, but is also roughly 4$\times$ slower, despite being implemented in the same conditions as the Neural CDE.
D.3 Dynamics Reconstruction
@@
As mentioned in LABEL:fig:readout_matrix, WARP uses a diagonal readout matrix $\theta_{t}(\tau)$ to self-decode the hidden states. This implies that post training, the weights $\theta_{t}$ and the time $\tau=t/(T-1)$ should be correlated. We confirm this hypothesis by plotting the correlation coefficient between the vector $\theta_{0:T}$ and all time points across all samples in the test set. We observe a strong either positive or negative correlation between the two quantities (see Figure 10).
@@
Figure 10: Correlation between the root network’s weights $\theta_{t}$ and the time $\tau$ on the MSD problem; indicating strong linear dependence between the two.
D.4 Extended Ablation Insights
We provide quantitative results for the ablation studies described in Section 3.4. Figures and Tables in this section are captioned with the corresponding paragraph title in Section 3.4.
@@ D.4 Extended Ablation Insights
Figure 11: Eliminating the root network — Test-set MSEs on the SINE problem. The omission of $\theta_{t}$ in favor of directly fitting $\varphi$ (which we call WARP-Phys) results in a catastrophic degradation in model expressivity. Performance is almost as bad as the Neural ODE analyzed in Figure 5.
@@
Figure 15: Ablation of the reparameterization trick — On the electricity problems, if stochastic sampling during training is not used, the model only predicts the mean of the distribution, thereby ignoring high-frequency components or noise in the signal. Here this is illustrated with a prediction on the ETTm1 test split.
Appendix E Visualizations
Figure 16: Completed images from the MNIST test set using WARP. The same set of images is shown across three settings: (Top) $L=100$, (Middle) $L=300$, (Bottom) $L=600$. Along the columns, we whow 4 groups of results, each with Ground Truth (GT), Reconstruction (Recons), and Uncertainty, resulting in 12 total columns. As our model sees more steps, its forecasting improves and its uncertainty decreases.
@@ Appendix E Visualizations
Figure 19: Completed images using WARP on the CelebA test set at high-resolution, illustrating the failure mode of limitation with long-range dependencies, e.g. when $T=64\times 64=4,096$.
@@
Figure 21: Completed sequences from the Lotka test set using WARP.

### v2

@@ D.6 Dynamical System Reconstruction
The LV experiment is performed for 1500 epochs with a batch size of 1024. Training is conducted with a teacher forcing probability $p_{\text{forcing}}=1.0$, meaning the model is always fed the ground truth inputs during training. This is because this is a memorisation task, and the goal is for the model to predict the next token knowing the previous one. LSTM and GRU use the same hyperparameters, except with hidden states of sizes 2450 and 2850 respectively (see Table 9).
@@
D.7 Time Series Classification
For time series classification tasks, encompassing both the UEA datasets and the Spirals dataset, models are consistently trained in the non-AR mode, with the categorical cross-entropy loss. Across all these classification experiments, root weight are evolved without weight clipping, and no dynamic tanh activation is applied to their final outputs. Key training hyperparameters exhibit some variation across these diverse datasets: the learning rate is $10^{-5}$ for the Spirals dataset and most UEA datasets (e.g., Ethanol, Heartbeat, Motor, SCP1, SCP2), with the Worms dataset being an exception at $10^{-6}$. The number of training epochs varies widely, ranging from 800 for the Worms dataset, 4000 for Spirals, up to 6500 for the Ethanol UEA dataset, with other UEA datasets generally trained for several thousand epochs. Given our limitation of 16GB available VRAM memory, batch sizes also differ significantly; for instance, the Worms dataset uses a batch size of 40, other UEA datasets use batch sizes typically in the hundreds (from approximately 280 to 560), and the Spirals dataset employs a large batch size of 10000. Regarding data preprocessing, input data normalisation is applied for several UEA datasets (specifically Ethanol, Heartbeat, SCP1, and SCP2), but it is not used for others like EigenWorms and MotorImagery, nor is it required for the Spirals dataset.
This task uses positional encoding in addition to normalised time. The dimension $d$ and the denominator constant $C$ of the positional encoding defined in Eq. 5 [89] and used in concatenation with the normalised time on the UEA dataset, are presented in Table 10.
Table 10: Hyperparameters for positional encoding on the UEA datasets.
Worms SCP1 SCP2 Ethanol Heartbeat Motor
Dimension $d$ 20 10 10 10 10 10
Denominator constant $C$ 20 10 10 10 5 10
Appendix E Additional Results
E.1 2D Image Experiments
@@ Appendix E Additional Results
E.2 Spirals & Neural CDEs
[…]
E.3 Computational Efficiency Comparison
To provide a comprehensive analysis of computational efficiency, we evaluate several key performance metrics for WARP and our baseline models. The experiments were conducted on an NVIDIA RTX 4080 GPU, ensuring a consistent hardware environment for all comparisons.
For the MNIST image completion task, we report the average wall-clock training time per epoch, peak GPU memory usage, and total parameter counts. To ensure a fair comparison, all models were trained with a fixed batch size of 128. The results, presented in Table 13, demonstrate WARP’s notable efficiency. Despite having a comparable number of parameters to the Transformer model, WARP requires significantly less GPU memory—on par with the much simpler GRU and LSTM architectures—and achieves the fastest training time.
Table 13: Training efficiency comparison on the MNIST image completion task. We report the average wall-clock time per epoch, peak GPU usage, and the number of learnable parameters. WARP is the most efficient in terms of both time and memory.
Model
Avg. training time
per epoch (seconds)
Peak GPU usage
Parameters
GRU 57.04 4.49 1.69
LSTM 59.22 4.95 1.70
S4 61.53 12.60 1.71
Transformer 18.62 10.03 1.69
WARP 45.22 2.89 1.69
A similar analysis was conducted for the UEA benchmark datasets, with results detailed in Table 14. For these experiments, the batch size was fixed to 32 across all models. The table provides a detailed breakdown of the training time, memory usage, and model complexity for WARP on each dataset.
Table 14: Detailed training metrics for WARP on the UEA benchmark datasets.
Dataset
Training time
per epoch (s)
Peak GPU
usage (MiB)
Num. of
epochs
Training batches
per epoch
Parameters
Worms 10.29 14598 1000 6 5.697
SCP1 0.92 654 5000 13 0.476
SCP2 3.85 2866 5000 9 17.34
Ethanol 2.22 1536 6500 12 4.681
Heartbeat 6.50 4354 1500 9 75.02
Motor 2.46 2558 2000 9 4.503
It is important to note that our implementation of WARP, along with the GRU, LSTM, and S4 baselines, utilises JAX. In contrast, the Transformer and all other baselines is implemented in PyTorch. This difference in framework can influence performance measurements. Following standard practice, we exclude any one-time JIT-compilation costs from the reported wall-clock times.
E.4 Normalised Time Correlation on Dynamics Reconstruction
@@
Let’s analyse when the root network takes as input exclusively the normalised time. In that case, WARP uses a diagonal readout matrix $\theta_{t}(\tau)$ as seen in LABEL:fig:readout_matrix to self-decode the hidden states. This implies that post training, the weights $\theta_{t}$ and the time $\tau=t/(T-1)$ should be correlated. We confirm this hypothesis by plotting the correlation coefficient between the vector $\theta_{0:T}$ and all time points across all samples in the test set. We observe a strong either positive or negative correlation between the two quantities (see LABEL:fig:correlation).
@@
(a) (a) Example “readout” matrix on the MSD problem for all time steps $t$ at all times $\tau$, highlighting WARP’s diagonal decoding direction $\theta_{t}(\tau)$; (b) Correlation between the root network’s weights $\theta_{t}$ and the time $\tau$ on the MSD problem; indicating strong linear dependence between the two.
E.5 Ablation Studies
We briefly discuss several experiments carried out to gain insights into our model. For all ablation studies, experimental protocols like training hyperparameters are presented in Appendix D. Figures and Tables in this section are captioned with the corresponding paragraph title.
Eliminating the root network.
The root network $\theta_{t}$ is integral to the efficacy of WARP. Although not an absolute prerequisite for the WARP-Phys variant, it nonetheless persists as a pivotal constituent of our framework. Illustratively, the omission of $\theta_{t}$ in favour of directly fitting $\varphi$ for the SINE modelling problem results in a catastrophic degradation in model expressivity.
@@ E.5 Ablation Studies
Figure 9: Eliminating the root network — Test-set MSEs on the SINE problem. The omission of $\theta_{t}$ in favor of directly fitting $\varphi$ (which we call WARP-Phys) results in a catastrophic degradation in model expressivity. Performance is almost as bad as the Neural ODE analysed in LABEL:fig:data_efficiency.
Initial network configuration.
Since WARP’s weight trajectory is driven by the changes in the signal and not the signal itself (see Eq. 1), it is important to have an expressive initial hypernetwork $\phi:\mathbf{x}_{0}\mapsto\theta_{0}$, which embeds the initial tokens into suitable weight spaces [51]. Our empirical investigations reveal that sidestepping this component substantially curtails the model performance on complex synthetic benchmarks, such as MSD-Zero, and on real-world datasets, including ETT.
@@
Data efficiency.
With $L=1$, the SINE benchmark is a challenging initial value problem. At equal (root) neural network parameter counts, we vary the number of training samples, and we plot MSE and MAPE test metrics for WARP and the Neural ODE [17]. The results, depicted in LABEL:fig:data_efficiency, not only show improved performance across data regimes, but they also indicate that more data is not necessarily better for WARP’s performance, suggesting potential for monotone learning [13].
(a) (a) Sample efficiency on SINE. (b) A dense weights-to-weights AA matrix on ETTm1.
Dense state transitions & Channel mixing.
@@
(a) Root network evaluation — When using normalised time as the coordinate system, if we fix the evaluation point $\tau$, we observe mild degradation in the qualitative results. While these figures are shown for MNIST with $L=300$, the behaviour is observed across problems, including dynamical systems like MSD (see LABEL:fig:readout_matrix). GT stands for the Ground Truth, Recons is for the Reconstruction/Completion, and Uncertainty is the model-outputted standard deviation.
Table 16: Positional Encodings (PE) ablation — We report the classification accuracy (%) of WARP with and without PE on the UEA datasets. The results show a consistent performance drop when PE is removed, underscoring its importance for long-range dependencies such as Worms and Motor.
Worms SCP1 SCP2 Ethanol Heartbeat Motor
with PE 70.93 $\pm$ 2.7 83.53 $\pm$ 2.0 57.89 $\pm$ 1.4 32.91 $\pm$ 4.2 88.65 $\pm$ 1.9 56.14 $\pm$ 5.1
w/o PE 60.98 $\pm$ 3.1 80.00 $\pm$ 2.0 57.89 $\pm$ 1.5 31.65 $\pm$ 0.8 77.42 $\pm$ 2.2 50.88 $\pm$ 2.3
Figure 13: Ablation of the reparametrisation trick — On the electricity problems, if stochastic sampling during training is not used, the model only predicts the mean of the distribution, thereby ignoring high-frequency components or noise in the signal. Here, this is illustrated with a prediction on the ETTm1 test split.
Appendix F Visualisations
Figure 14: Completed images from the MNIST test set using WARP. The same set of images is shown across three settings: (Top) $L=100$, (Middle) $L=300$, (Bottom) $L=600$. Along the columns, we show 4 groups of results, each with Ground Truth (GT), Reconstruction (Recons), and Uncertainty, resulting in 12 total columns. As our model sees more steps, its forecasting improves and its uncertainty decreases.
@@ Appendix F Visualisations
Figure 17: Completed images on the CelebA test set at high-resolution ($T=64\times 64=4096$), using positional encoding [89]. This illustrates WARP’s suitability for long-range dependencies.
@@
Figure 19: Completed sequences from the LV test set using WARP.